# Michel F. C. Haddad

### v1

@@ 4 Estimation
$-\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}.$ (23)
A cross-fitted estimator of the variance in equation (21) may be constructed in an analogous manner as outlined for the case of repeated cross-sections. Lastly, the corresponding variance expressions for the estimation of $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$, the ATET of a treatment dose in an earlier period on an outcome in a later period as in given (11), is obtained by a appropriately defining treatment and outcome periods according to the time lags $t-s$.
As an alternative approach to the asymptotic variance approximation, one may apply a multiplier bootstrap procedure for constructing confidence intervals. Let $\{\xi_{i}\}_{i=1}^{n}$ be an i.i.d. sequence of sub-exponential random variables independent from the data $\{W_{i}\}_{i=1}^{n}$, such that $E[\xi]=Var(\xi)=1$. In each bootstrap replication $b=1,\cdots,B$, one draws such a sequence $\{\xi_{i}\}_{i=1}^{n}$ and then constructs a bootstrap-specific ATET estimators, denoted by $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}$, using the following expression:
$\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\cdot\psi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta}).$ (24)
Denoting by $\hat{c}_{\alpha}$ the $\alpha$th quantile of the difference $\{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}-\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}\}_{b=1}^{B}$, a $1-\alpha$ confidence interval may be constructed based on $[\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\hat{c}_{1-\alpha/2},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\hat{c}_{\alpha/2}]$.
@@ 5 Asymptotic Theory
This section establishes the asymptotic properties of our estimators. The results largely follow Zhang (2024) and we sketch the main idea of the proof in Appendix A.2. First, we impose a set of regularity conditions on the kernel function and the boundedness and smoothness of the model parameters. These conditions are used to establish three main results. First, the bias of ATET from using a kernel approximation becomes negligible asymptotically if the kernel bandwidth is undersmoothed. Second, the ATET estimators are asymptotically normal. Third, the variance estimators based on the asymptotic variance formulas proposed in Section 4 are consistent. We focus on the case without time lags, and the results for the case with time lags follow immediately by appropriately defining treatment and outcome periods according to the time lag $t-s$. Recalling that the weighting function is defined as $\omega(D;d,h)=\frac{1}{h}\mathcal{K}\left(\frac{D-d}{h}\right)$ with $h$ being the bandwidth, we consider bandwidth-dependent versions of the generalized propensitiy scores, denoted by $\rho_{d,t}^{h}(\mathbf{D_{t-1}},X)=E[\omega(D;d,h)\cdot I\{T=t\}|\mathbf{D_{t-1}},X]$ and $p_{d}^{h}(\mathbf{D_{t-1}},X)=E[\omega(D;d,h)|\mathbf{D_{t-1}},X]$, in some the subsequent assumptions. Furthermore, we denote by $\hat{\rho}_{d,t}^{h}(\mathbf{D_{t-1}},X)$ and $\hat{p}_{d}^{h}(\mathbf{D_{t-1}},X)$ estimates of $\rho_{d,t}^{h}(\mathbf{D_{t-1}},X)$ and $p_{d}^{h}(\mathbf{D_{t-1}},X)$.
@@
[…]
@@
(a) For some constants $10<c<$ 1 and $0<C<\infty$, $\Pi_{d_{t},t}>c$, $|\rho_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)|>c$, $|\rho_{d_{t}^{\prime},t}^{h}(\mathbf{D_{t-1}},X)|>c$, $|\rho_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)|>c$, $|Y_{T}|<C$, $|\rho_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)|<C$, $|\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)|<C$, $|\mu_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|<C$, and $|\mu_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)|<C$ almost surely; (b) $\Pi_{d_{t},t}\in C^{2}(\text{supp}(D_{t}))$ and $|\Pi_{d_{t},t}^{(2)}|<\infty$; (c) $\rho_{d_{t},t-1}(\mathbf{D_{t-1}},X)\in C^{2}(\text{supp}(D_{t}))$ and $\sup_{x\in\text{supp}(X)}|\rho_{d_{t},t-1}^{(2)}(\mathbf{D_{t-1}},X)|<\infty$; (d) joint density $f_{Y_{T},D_{t}}(y,d_{t})\in C^{2}(\text{supp}(Y_{T}))$ and $\sup_{d_{t}\in\text{supp}(D_{t})}|f_{Y_{T},D_{t}}^{(2)}(y,d_{t})|<\infty$.
@@
(a) The kernel bandwidth $h$ is a deterministic sequence that depends on $n$ and satisfies $nh\to\infty$ and $\sqrt{nh^{5}}=o(1)$; (b) there exists a sequence $\varepsilon_{n}\to 0$, such that $h^{-1}\varepsilon_{n}^{2}=o(1)$; (c) with probability tending to $1$, $\|\hat{\rho}_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)-\rho_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq h^{-1/2}\varepsilon_{n}$, $\|\hat{\rho}_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)-\rho_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\rho}_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)-\rho_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}}(t-1,\mathbf{D_{t-1}},X)-\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)-\mu_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-\mu_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$; (d) with probability tending to $1$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{\rho}_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)>c$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{\rho}_{d_{t}^{\prime},t}^{h}(\mathbf{D_{t-1}},X)>c$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{\rho}_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)>c$, $\|\hat{\rho}_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, and $\|\hat{\mu}_{d_{t}}(t-1,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, $\|\hat{\mu}_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, $\|\hat{\mu}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$.
@@
(a) For some constants $10<c<$ 1 and $0<C<\infty$, $P_{d_{t}}>c$, $|Y_{t-1}|<C$, $|Y_{t}|<C$, $p_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)>c$, $|p_{d_{t}}^{h}(\mathbf{D_{t-1}},X)|<C$, and $|m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|<C$ almost surely for all $h>0$; (b) $P_{d_{t}}\in C^{2}(\text{supp}(D_{t}))$ and $|P_{d_{t}}^{(2)}|<\infty$; (c) $p_{d_{t}}(\mathbf{D_{t-1}},X)\in C^{2}(\text{supp}(D_{t}))$ and $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}|p_{d_{t}}^{(2)}(\mathbf{D_{t-1}},X)|<\infty$; (d) for $\Delta Y=Y_{t}-Y_{t-1}$, the joint density $f_{\Delta Y,D_{t}}(s,d_{t})\in C^{2}(\text{supp}(\Delta Y))$ and $\sup_{d\in\text{supp}(D_{t})}|f_{\Delta Y,D_{t}}^{(2)}(s,d_{t})|<\infty$.
@@
(a) The kernel bandwidth $h$ is a deterministic sequence that depends on $n$ and satisfies $nh\to\infty$ and $\sqrt{nh^{5}}=o(1)$; (b) there exists a sequence $\varepsilon_{n}\to 0$ such that $h^{-1}\varepsilon_{n}^{2}=o(1)$; (c) with probability tending to $1$, $\|\hat{p}_{d_{t}}^{h}(\mathbf{D_{t-1}},X)-p_{d_{t}}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq h^{-1/2}\varepsilon_{n}$, $\|\hat{p}_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)-p_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{m}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$; (d) with probability tending to $1$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{p}_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)>c$, $\|\hat{p}_{d_{t}}^{h}(\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, $\|\hat{m}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$.
Assumption 9 concerning the kernel function is standard and many commonly used smoothing functions satisfy these conditions, like the Gaussian or Epanechnikov kernel. Furthermore, we impose two sets of regularity conditions in both repeated cross-sections and panel settings. Firstly, we require that bandwidth-dependent nuisance parameters in the population are bounded and satisfy certain smoothness conditions, see Assumptions 10 and 12. Secondly, we impose high-level assumptions on the quality of the estimators of the nuisance parameters, permitting those estimators to converge at a slower rate compared to a setting without Neyman orthogonality, see Assumptions 11 and 13. A first result obtained from these conditions is that the biases in repeated cross-sections and panel settings are of order $O(h^{2})$, such that they vanish asymptotically when using an undersmoothed kernel bandwidth for ATET estimation.
@@
Suppose Assumptions 9, 10, 11 hold for the repeated cross-sections case, and Assumptions 9, 12, 13 hold for the panel case. Then, the estimation bias is given by $B(h)=\Delta_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ for $d_{t},d_{t}^{\prime}\in\text{supp}(D_{t})$.
The next two theorems are the main results of this section. The first theorem states that our proposed ATET estimators are asymptotically normal and converge at $\sqrt{nh}$ rate to the true ATET, while the second theorem states that our variance estimators proposed in the previous section are consistent.
@@
Suppose Assumptions 4, 5, 6, 7 hold. Moreover, suppose Assumptions 9, 10, 11 hold for the repeated cross-sections case, and Assumptions 9, 12, 13 hold for the panel case. Then the ATET estimators defined in Section 4 satisfy
@@
where $\sigma_{h}$ is defined as in (16) and (21) for the repeated cross-sections case and panel case, respectively.
@@
Assume Theorem 1 holds and assume that $h^{-2}\varepsilon_{n}^{2}=o(1)$. Then the variance estimators defined in Section 4 are consistent for the asymptotic variance $\sigma_{h}$ defined in (16) and (21) for the repeated cross-sections case and panel case, respectively.

### v2

@@ 5 Estimation
$-\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D}_{t-1},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)]}{p_{d_{t}^{\prime}}(\mathbf{D}_{t-1},X)\cdot P_{d_{t}}}.$
A cross-fitted estimator of the variance in equation (27) can be constructed analogously as outlined for the case of repeated cross-sections.
Lastly, the corresponding estimators and variance expressions for the ATET with lagged treatment $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$ can be easily obtained by substituting the expressions shown earlier with the lagged versions.
As an alternative approach to the asymptotic variance approximation, one can employ a multiplier bootstrap procedure for constructing confidence intervals. Let $\{\xi_{i}\}_{i=1}^{n}$ be an i.i.d. sequence of sub-exponential random variables independent from the data $\{W_{i}\}_{i=1}^{n}$, such that $E[\xi]=Var(\xi)=1$. In each bootstrap replication $b=1,\cdots,B$, one draws such a sequence $\{\xi_{i}\}_{i=1}^{n}$ and then constructs a bootstrap-specific ATET estimators, denoted by $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}$, using the following expression:
[…]
Let $\hat{c}_{\alpha}$ denote the $\alpha$-th quantile of the difference $\{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}-\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}\}_{b=1}^{B}$, a $1-\alpha$ confidence interval may be constructed as $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\hat{c}_{1-\alpha/2},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\hat{c}_{\alpha/2}]$.
Moreover, in the continuous treatment setting, it is also useful to consider uniform inference, and the bootstrap estimator defined in (29) can be used to construct the uniform confidence bands. For simplicity, we focus on the two-period repeated cross-sectional case comparing the treated group with intensity $d$ and the control group with intensity 0. In particular, this extends the repeated cross-sectional results in Zhang (2025) to allow time-varying covariates using efficient scores.
Recall that our causal parameter in this setting is defined as
$\Delta_{d_{t},0,t}=E[Y_{t}(d_{t})-Y_{t}(0)|D_{t}=d_{t},T=t],$ (30)
and we aim to establish valid uniform confidence bands for $\Delta_{d_{t},0,t}$ over the support of $D_{t}$. Let $\hat{\Delta}_{d_{t},0,t}^{h}$, $\hat{\Delta}_{d_{t},0,t}^{b,h}$, and $\hat{\sigma}_{h}^{2}(d_{t},0,t)$ denote our DML estimator, the multiplier bootstrap estimator, and the cross-fitted variance estimator respectively. We consider the following procedure that follows Chernozhukov et al. (2014b, a), Fan et al. (2022), and Zhang (2025) to establish valid uniform confidence bands.
  • 1.
    Estimate $\hat{\Delta}_{d_{t},0,t}^{h}$ and $\hat{\sigma}_{h}^{2}(d_{t},0,t)$ on a finite grid of values $d_{t}\in\bar{\mathcal{D}}_{t}\subset supp(D_{t})$.
  • 2.
    For each $b=1,\cdots,B$, draw an i.i.d. sequence of multipliers $\{\xi\}_{i=1}^{N}$ from a $N(1,1)$ distribution, and construct $\hat{\Delta}_{d_{t},0,t}^{b,h}$ for all $d_{t}\in\bar{\mathcal{D}}_{t}$.
  • 3.
    Compute $\hat{c}(1-\alpha)$, which we denote as the $1-\alpha$-quantile of
    $\Bigg\{\max_{d_{t}\in\bar{\mathcal{D}}_{t}}\frac{\sqrt{N}|\hat{\Delta}_{d_{t},0,t}^{h}-\hat{\Delta}_{d_{t},0,t}^{b,h}|}{\hat{\sigma}_{h}^{2}(d_{t},0,t)}\Bigg\}_{b=1}^{B}.$
  • 4.
    For all $d_{t}\in supp(D_{t})$, construct the $1-\alpha$ uniform confidence band as
    $\hat{\Delta}_{d_{t},0,t}^{h}-\hat{c}(1-\alpha)\hat{\sigma}_{h}^{2}(d_{t},0,t)/\sqrt{N},\quad\hat{\Delta}_{d_{t},0,t}^{h}+\hat{c}(1-\alpha)\hat{\sigma}_{h}^{2}(d_{t},0,t)/\sqrt{N}].$
We omit the formal theoretical discussion here and point the readers to Zhang (2025), which establishes valid uniform asymptotic linear expansion of the bootstrap estimators analogous to ours, and the validity of the uniform confidence bands constructed above follows from the results in Chernozhukov et al. (2014b, a) and Fan et al. (2022).
@@ 6 Asymptotic Theory
This section establishes the asymptotic properties of our estimators. The results largely follow Zhang (2025) and we sketch the main idea of the proofs in Appendix A.2. First, we impose a set of regularity conditions on the kernel function and the boundedness and smoothness of the model parameters. These conditions are used to establish three main results. First, the bias of ATET from using a kernel approximation becomes negligible asymptotically if the kernel bandwidth is undersmoothed. Second, the ATET estimators are asymptotically normal. Third, the variance estimators based on the asymptotic variance formulas proposed in Section 5 are consistent. We focus on the case without time lags, and the results for the case with time lags follow immediately by appropriately defining treatment and outcome periods according to the time lag tst-s. Recall that in the previous sections, we define the weighting function as $\omega(D;d,h)=\frac{1}{h}\mathcal{K}\left(\frac{D-d}{h}\right)$ with $h$ being the bandwidth. Using this weighting function, we consider the bandwidth-dependent approximations of the generalized propensitiy scores, denoted by $\rho_{d,t}^{h}(\mathbf{D}_{t-1},X)=E[\omega(D;d,h)\cdot I\{T=t\}|\mathbf{D}_{t-1},X]$ and $p_{d}^{h}(\mathbf{D}_{t-1},X)=E[\omega(D;d,h)|\mathbf{D}_{t-1},X]$, for the cross-sectional and panel settings respectively. In addition, we use $\hat{\rho}_{d,t}^{h}(\mathbf{D}_{t-1},X)$ and $\hat{p}_{d}^{h}(\mathbf{D}_{t-1},X)$ to denote the estimates of $\rho_{d,t}^{h}(\mathbf{D}_{t-1},X)$ and $p_{d}^{h}(\mathbf{D}_{t-1},X)$.
@@
The kernel function $\mathcal{K}(\cdot)$ satisfies: (a) $\mathcal{K}(\cdot)$ is bounded and differentiable; (b) $\int\mathcal{K}(u)du=1$, $\int u\mathcal{K}(u)du=0$, $0<\int u^{2}\mathcal{K}(u)du<\infty$.
@@
(a) There exists $0<c<1$ and $0<C<\infty$ such that $\Pi_{d_{t},t}>c$, $c<\rho_{d_{t},t-1}^{h}(\mathbf{D}_{t-1},X)<C$, $\rho_{d_{t}^{\prime},t}^{h}(\mathbf{D}_{t-1},X)>c$, $\rho_{d_{t}^{\prime},t-1}^{h}(\mathbf{D}_{t-1},X)>c$, $|Y_{T}|<C$, $|\mu_{d_{t}}(t-1,\mathbf{D}_{t-1},X)|<C$, $|\mu_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)|<C$, and $|\mu_{d_{t}^{\prime}}(t-1,\mathbf{D}_{t-1},X)|<C$ almost surely; (b) $\Pi_{d_{t},t}\in C^{2}(\text{supp}(D_{t}))$ and $|\Pi_{d_{t},t}^{(2)}|<\infty$; (c) $\rho_{d_{t},t-1}(\mathbf{D}_{t-1},d)\in C^{2}(\text{supp}(D_{t}))$ for all $(\mathbf{D}_{t-1},x)$, and $\sup_{d_{t},d_{t-1},x}|\rho_{d_{t},t-1}^{(2)}(\mathbf{D}_{t-1},x)|<\infty$; (d) joint density $f_{Y_{T},D_{t}}(y,d_{t})\in C^{2}(\text{supp}(Y_{T}))$ and $\sup_{y,d_{t}\in\text{supp}(Y_{T},D_{t})}|f_{Y_{T},D_{t}}^{(2)}(y,d_{t})|<\infty$.
@@
(a) The kernel bandwidth $h=h_{N}\to 0$ satisfies $Nh\to\infty$ and $\sqrt{Nh^{5}}=o(1)$; (b) there exists a sequence $\varepsilon_{n}\to 0$, such that $h^{-1}\varepsilon_{n}^{2}=o(1)$; (c) with probability tending to $1$, $\|\hat{\rho}_{d_{t},t}^{h}(\mathbf{D}_{t-1},X)-\rho_{d_{t},t}^{h}(\mathbf{D}_{t-1},X)\|_{P,2}\leq h^{-1/2}\varepsilon_{n}$, $\|\hat{\rho}_{d_{t},t-1}^{h}(\mathbf{D}_{t-1},X)-\rho_{d_{t},t-1}^{h}(\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\rho}_{d_{t}^{\prime},t-1}^{h}(\mathbf{D}_{t-1},X)-\rho_{d_{t}^{\prime},t-1}^{h}(\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}}(t-1,\mathbf{D}_{t-1},X)-\mu_{d_{t}}(t-1,\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}^{\prime}}(t-1,\mathbf{D}_{t-1},X)-\mu_{d_{t}^{\prime}}(t-1,\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)-\mu_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$; (d) with probability tending to $1$, $c<\hat{\rho}_{d_{t},t-1}^{h}(\mathbf{D}_{t-1},X)<C$, $\hat{\rho}_{d_{t}^{\prime},t}^{h}(\mathbf{D}_{t-1},X)>c$, and $\hat{\rho}_{d_{t}^{\prime},t-1}^{h}(\mathbf{D}_{t-1},X)>c$ almost surely; (e) with probability tending to 1, $\|\hat{\mu}_{d_{t}}(t-1,\mathbf{D}_{t-1},X)\|_{P,\infty}<C$, $\|\hat{\mu}_{d_{t}^{\prime}}(t-1,\mathbf{D}_{t-1},X)\|_{P,\infty}<C$, and $\|\hat{\mu}_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)\|_{P,\infty}<C$.
@@
(a) There exist constants $0<c<1$ and $0<C<\infty$, $P_{d_{t}}>c$, $|Y_{t-1}|<C$, $|Y_{t}|<C$, $p_{d_{t}^{\prime}}^{h}(\mathbf{D}_{t-1},X)>c$, $|p_{d_{t}}^{h}(\mathbf{D}_{t-1},X)|<C$, and $|m_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)|<C$ almost surely; (b) $P_{d_{t}}\in C^{2}(\text{supp}(D_{t}))$ and $|P_{d_{t}}^{(2)}|<\infty$; (c) $p_{d_{t}}(\mathbf{D}_{t-1},x)\in C^{2}(\text{supp}(D_{t}))$ for all $(\mathbf{D}_{t-1},x)$ and $\sup_{\text{supp}(\mathbf{D}_{t-1},X)}|p_{d_{t}}^{(2)}(\mathbf{D}_{t-1},X)|<\infty$; (d) for $\Delta Y=Y_{t}-Y_{t-1}$, the joint density $f_{\Delta Y,D_{t}}(s,d_{t})\in C^{2}(\text{supp}(\Delta Y))$ and $\sup_{s,d_{t}\in\text{supp}(\Delta Y,D_{t})}|f_{\Delta Y,D_{t}}^{(2)}(s,d_{t})|<\infty$.
@@
(a) The kernel bandwidth $h=h_{N}\to 0$ satisfies $Nh\to\infty$ and $\sqrt{Nh^{5}}=o(1)$; (b) there exists a sequence $\varepsilon_{n}\to 0$ such that $h^{-1}\varepsilon_{n}^{2}=o(1)$; (c) with probability tending to $1$, $\|\hat{p}_{d_{t}}^{h}(\mathbf{D}_{t-1},X)-p_{d_{t}}^{h}(\mathbf{D}_{t-1},X)\|_{P,2}\leq h^{-1/2}\varepsilon_{n}$, $\|\hat{p}_{d_{t}^{\prime}}^{h}(\mathbf{D}_{t-1},X)-p_{d_{t}^{\prime}}^{h}(\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{m}_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)-m_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)\|_{P,2}\leq\varepsilon_{n}$; (d) with probability tending to $1$, $\sup_{\text{supp}(\mathbf{D}_{t-1},X)}\hat{p}_{d_{t}^{\prime}}^{h}(\mathbf{D}_{t-1},X)>c$, $\|\hat{p}_{d_{t}}^{h}(\mathbf{D}_{t-1},X)\|_{P,\infty}<C$, $\|\hat{m}_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)\|_{P,\infty}<C$.
Assumption 11 on the kernel function is standard and many commonly used smoothing functions satisfy these conditions, e.g., the Gaussian or Epanechnikov kernels. Moreover, we impose two sets of regularity conditions in both repeated cross-sections and panel settings. First, we require that bandwidth-dependent nuisance parameters in the population are bounded and satisfy certain smoothness conditions, see Assumptions 12 and 14. Second, we impose high-level assumptions on the quality of the estimators of the nuisance parameters, permitting those estimators to converge at a slower rate compared to a setting without Neyman orthogonality, see Assumptions 13 and 15. A first result obtained from these conditions is that the biases in repeated cross-sections and panel settings are of order $O(h^{2})$, such that they vanish asymptotically when using an undersmoothed kernel bandwidth for ATET estimation.
@@
Suppose Assumptions 11, 12, 13 hold for the repeated cross-sections case, and Assumptions 11, 14, 15 hold for the panel case. Then, the bias satisfies $B(h)=\Delta_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}=O(h^{2})$ for $d_{t},d_{t}^{\prime}\in\text{supp}(D_{t})$.
The next two theorems are the main results of this section. The first theorem states that our proposed ATET estimators are asymptotically normal and converge at $\sqrt{nh}$ rate to the true ATET, while the second theorem states that our cross-fitted variance estimators proposed in the previous section are consistent.
@@
Suppose Assumptions 4, 5, 6, 7 hold. Moreover, suppose Assumptions 11, 12, 13 hold for the repeated cross-sections case, and Assumptions 11, 14, 15 hold for the panel case. Then for $\varepsilon_{n}=o(n^{-1/4})$, the ATET estimators defined in Section 5 satisfy
@@
where $\sigma_{h}$ is defined as in (25) and (27) for the repeated cross-sections case and panel case respectively.
@@
Assume the conditions in Theorem 1 hold and assume that $h^{-2}\varepsilon_{n}^{2}+h^{-3}n^{-1}=o(1)$. Then the cross-fitted variance estimators defined in Section 5 are consistent for the asymptotic variance $\sigma_{h}$ defined in (25) and (27) for the repeated cross-sections case and panel case respectively.